OpenStudy (anonymous):

In this reaction had an 85.0% yield, how much CH4 was produced from 1250g C

OpenStudy (aaronq):

start by writing a balanced equation

OpenStudy (anonymous):

but what is my equation going to contain if I have one compound?

OpenStudy (anonymous):

@aaronq

OpenStudy (anonymous):

oh wait 2C + 2H20 ---> CH4 + CO2 okay know what

OpenStudy (aaronq):

hmm i would've done \(C + 2H_2 \rightarrow CH_4\)

OpenStudy (aaronq):

so now you wanna convert the mass of C to moles, then find the moles of \(CH_4\) that can be produced with that much C. Use the stoichiometric coefficients to find moles produced. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) Now, since they said you have an 85% yield, multiply the moles of CH4 you found by 0.85 to finish the question.